OpenStudy (anonymous):

help with systems of equations ! solve x + 2y = 13 4x + 8y = -9

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (nikato):

There are to ways to solve. Substitution or elimination. Which way do u want to use?

OpenStudy (anonymous):

elimination!!

OpenStudy (anonymous):

wait a sec is it 3,5 ?

OpenStudy (nikato):

No, that's not correct. So for elimination, the first thing I want to do is multiply 4 on both sides in the first equation

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

4x + 8y = 52

OpenStudy (anonymous):

and when you add you get 8x + 16y = 43

OpenStudy (nikato):

Then u want to subtract the 2 equations

OpenStudy (anonymous):

isn't it add

OpenStudy (nikato):

No, u subtract. If u want to add, u can multiply one of the equaions by -1. Whether u add or subtract, ur goal is to try to elimate one of the variables

OpenStudy (anonymous):

ok.. so i got 61

OpenStudy (anonymous):

actually 0x + 0y = 61

OpenStudy (nikato):

Correct. Which is also 0=61. And since this isn't true, then this system of equaions has no solutions which means that they are parallel lines

OpenStudy (anonymous):

so the answer is no solutions.

OpenStudy (nikato):

Correct. Because 0 doesn't not equal 61

OpenStudy (anonymous):

ok thanks

OpenStudy (nikato):

No problem

OpenStudy (anonymous):

i have a couple more i don't need you to do the whole thing because i know its long but i just need to know what to multiply the first equation by ,

OpenStudy (anonymous):

like this one 3x - 5y = 17 9x - 15y = -4

OpenStudy (anonymous):

do i multiply 9 to both sides of the first one ?

OpenStudy (nikato):

Ok. So u can look at the x's. one has a 3 and the other has a 9. In order to get rid of them, u would have to make them =. Since 3x3=9, u can just multiply the first equation both sides by 3

OpenStudy (anonymous):

ahh so the goal is to get opposite coefficients

OpenStudy (anonymous):

then we add or subtract ?

OpenStudy (nikato):

Becuz after multiplying, u have 9x and 9x and in order to make it=0 u would have to subtract. So in this case, u would subtract

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

ok i think im stuck i got -30y = 13

OpenStudy (nikato):

I think u did something wrong 9x-15y=51 - (9x-15y=-4) ------------------ 0x+0y=55 So 0=55

OpenStudy (anonymous):

wow.

OpenStudy (anonymous):

you're so smart lol

OpenStudy (anonymous):

so its no solution again ?

OpenStudy (nikato):

Haha. Thanks. And yes, no solution is correct

OpenStudy (anonymous):

are these hard for you

OpenStudy (nikato):

Not really especially compared to what I'm doing now

OpenStudy (anonymous):

calculus

OpenStudy (nikato):

No, geometry h

OpenStudy (anonymous):

oh ok